# Tensor Product Attention Is All You Need

### v5

@@ F.1 Multi-Query Attention (MQA)
Multi-Query Attention (MQA) [46] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection. Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
@@ F.2 Grouped Query Attention (GQA)
Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups. Within each group, heads share a common key and value projection, while each head maintains its own unique query projection. Formally, let $g(i)$ denote the group index for head $i\in\{1,\dots,h\}$, where $g(i)\in\{1,\dots,G\}$. The projections are:
@@ F.3 Multi-head Latent Attention (MLA)
Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [32] and DeepSeek-V3 [33], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
@@
Here, $\bm{W}^{DKV}\in\mathbb{R}^{d_{\text{model}}\times d_{c}}$ projects to a compressed dimension $d_{c}$, $\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed keys, $\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}$ projects to a residual key component for RoPE, and $\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed values. $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the shared compressed KV latent (where $d_{c}\ll d_{h}h$). The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [60].
@@
where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix. The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding. However, as noted by [51], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations. RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4). If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
@@ F.4 Multi-matrix Factorization Attention (MFA)
[21] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
@@ F.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; 52) to encode positional information in the query/key vectors. Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head). $\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$. The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$. If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
@@ Appendix G More on TPA
Parameter Initialization for TPA Factors. We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [15]. Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$. Here, $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrix. This initialization strategy is chosen to help maintain the variance of activations and gradients as they propagate through the network layers, contributing to stable training.
@@ H.1 Experimental Settings
Training Setup Details. We follow the nanoGPT training configuration [24]. In particular, we use the AdamW [35] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [36] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).

### v6

@@ F.1 Multi-Query Attention (MQA)
Multi-Query Attention (MQA) [shazeer2019fast] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection. Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
@@ F.2 Grouped Query Attention (GQA)
Grouped Query Attention (GQA) [ainslie2023gqa] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups. Within each group, heads share a common key and value projection, while each head maintains its own unique query projection. Formally, let $g(i)$ denote the group index for head $i\in\{1,\dots,h\}$, where $g(i)\in\{1,\dots,G\}$. The projections are:
@@ F.3 Multi-head Latent Attention (MLA)
Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [liu2024deepseek] and DeepSeek-V3 [liu2024deepseekv3], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
@@
Here, $\bm{W}^{DKV}\in\mathbb{R}^{d_{\text{model}}\times d_{c}}$ projects to a compressed dimension $d_{c}$, $\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed keys, $\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}$ projects to a residual key component for RoPE, and $\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed values. $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the shared compressed KV latent (where $d_{c}\ll d_{h}h$). The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [vaswani2017attention].
@@
where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix. The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding. However, as noted by [kexuefm-10091], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations. RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4). If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
@@ F.4 Multi-matrix Factorization Attention (MFA)
[hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
@@ F.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; su2024roformer) to encode positional information in the query/key vectors. Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head). $\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$. The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$. If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
@@ Appendix G More on TPA
Parameter Initialization for TPA Factors. We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [glorot2010understanding]. Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$. Here, $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrix. This initialization strategy is chosen to help maintain the variance of activations and gradients as they propagate through the network layers, contributing to stable training.
@@ H.1 Experimental Settings
Training Setup Details. We follow the nanoGPT training configuration [Karpathy2022]. In particular, we use the AdamW [loshchilov2017decoupled] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [loshchilov2016sgdr] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).